OpenStudy (ragingsquirrel):

***HELP****HELP*** The area of the base of a rectangular prism is 24 square inches, and its height is a whole number. Which of the following could not be its volume? 120 in3 144 in3 160 in3 168 in3

OpenStudy (anonymous):

You know the volume of a prism is lwh. The base is given. Which is lw. Set lw = 24. plug this back into the original equation lwh = v 24h = v now substitute in the given volumes and isolate the height fro each set. h = 120/24 = 5 h = 144/24 = 6 h = 160/24 = 6.66666 h = 168/24 = 7 The question states the height is a whole number which eliminates the third option as a possible height.

OpenStudy (anonymous):

V = base * height -> h = v/base Since h is whole number -> V is divisable to base => Only V = 160 is undivisable to 24!